A gun of mass 10kg released a bullet of mass 20gm at a speed of 100m/s. Find the gun's speed of recoil
1
Expert's answer
2019-04-17T11:14:39-0400

Following momentum conservation law,


"Mv = mu"

"v = \frac{mu}{M} = \frac{20 \cdot 100}{10\cdot 10^3}=0.2 \, \text{m/s}"